OpenStudy (anonymous):

solve: 0=cos(2x)+3cosx-1, 0

OpenStudy (anonymous):

cos(2x) = 1 - 2cos^2x -2cos^2x + 3cosx = 0 let cosx = u -2u^2 + 3u = 0 use quadratic formula to solve for u once u is solved, cosx = u

OpenStudy (anonymous):

there cosx = 1.5 or cosx = 0

OpenStudy (anonymous):

therefore*

OpenStudy (anonymous):

ohh thank you

OpenStudy (anonymous):

no prob